OpenStudy (anonymous):

Please Help! Find the exact value: csec 22pi/3

OpenStudy (anonymous):

22pi/3 = (21pi + pi)/3 = 7pi + pi/3

OpenStudy (anonymous):

ok so for this question we know csc = 1/sin alright. and we also know that 22pi/3 is 1320. So one whole circle is 360 degrees and if we keep going around to 3 circles we end up at 1080. which is still 0 if you think about it then from there do 1320 - 1080 = 240 so basically now your finding your exact value of 240. and if you do that you get sin 240 = -sqrt(3)/2 and this implies csc 240 = -2/sqrt(3) :)

OpenStudy (anonymous):

it's in third quadrant.

OpenStudy (anonymous):

I think hamza is right

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Thanx guys for the help!